OpenStudy (anonymous):

Find tan θ if cos θ = 12/13 and sin θ<0

OpenStudy (anonymous):

use the identities sin^2+cos^2=1 and tan=sin/cos I think there's a pythagorean triple with 2 sides 12 and 13. I forget the other side.

OpenStudy (kainui):

Remember SOH CAH TOA You're given cos theta = 12/13 which is the CAH which means cosine is adjacent over hypotenuse. See what I did there? You now can use the pythagorean theorem to find that other side, which is what sine and tangent are found using SOH and TOA. Good luck!

OpenStudy (anonymous):

use my way if you like algebra more. use kainui's way if you like geometry more.

OpenStudy (anonymous):

Tan theta=5/12 since cos theta=Adj/hyp

OpenStudy (anonymous):

@KIEMA I'm getting -5/12, is this correct?

OpenStudy (anonymous):

@Peter14 thank you very much!

OpenStudy (anonymous):

jaytag is right, sin theta is less than zero